A 500mm diameter pipeline discharges water from a large reservoir. Determine the maximum possible elevation of its summit at B for a discharge of 2. 15m3/s without the absolute pressure becoming less than 20 KN/m2. Take atmospheric pressure as 10^6

1
Expert's answer
2021-11-17T03:39:02-0500


Pabs = P(atm) + P(gauge)

=20×103

Pabs,B = 9810 = 2.038 m

Patm = 10.3m

Pgauge,B = 2.038 − 10.3 = −8.26m

Q=AV→V=Q= 2.15 =10.95m/s

"=A \u03c0^4\u00d70.52"

Applying Bernoulli’s equation between the points A and B.

PA VA2 PB VB2

ρg+2g+zA =ρg+2g+zB +∑hL


"z_C =6.11m"